OpenStudy (anonymous):

solve each quadratic equation. Show your work 4x^2 = 100

OpenStudy (photon336):

well first you need to get it into the form ax^{2}+bx+C so you would subtract 100 from both sides. \[4x^{2}-100 = 0 \]

OpenStudy (anonymous):

will 100 cancel out?

OpenStudy (photon336):

on the right side it will but not the left so we would now get \[4x^{2}-100 = 0 \]

OpenStudy (anonymous):

ohh okay i think i know what to do

OpenStudy (photon336):

okay so show me the next step

OpenStudy (anonymous):

multiply by 2?

OpenStudy (photon336):

so you would factor out a 4. \[4(x^{2}-25) = 0 \] then divide both sides by 4 Which would give you this \[x^{2}-25 = 0 \] now how do you solve this?

OpenStudy (anonymous):

add a 5?

OpenStudy (photon336):

watch this we add 25 to both sides. \[x^{2}+25-25 = 0+25 \] \[x^{2} = 25 \]

OpenStudy (photon336):

how would we find x @kimyy563d ?

rebeccaxhawaii (rebeccaxhawaii):

*raises hand* I KNOW I KNOW PICK ME PICK ME

OpenStudy (photon336):

haha okay continue

OpenStudy (anonymous):

im confused

rebeccaxhawaii (rebeccaxhawaii):